Two resistors connected in series have an equivalent resistance of 608 Ω. When they are connected in parallel, their equivalent resistance is 122 Ω. Find the resistance of each resistor.

Your response differs from the correct answer by more than 10%. Double check your calculations. Ω (small resistance)

Your response differs from the correct answer by more than 10%. Double check your calculations. Ω (large resistance)
Answer #1

Let them R1 and R2,

Here it is given that, series equivalent R1 + R2 = 608

also parallel equivalent R1R2/[R1+R2] = 122

or R1*(608-R1)/608 = 122

R1 = 169   Ω

b] R2 = 608 - R1    = 439 Ω